OpenStudy (ranya99):

A watermelon cannon fires a watermelon vertically up into the air at a velocity of +9.5m/s m/s, starting from an initial position 1.2 meters above the ground. When the watermelon reaches the peak of its flight, what is: (a) its velocity (b) its acceleration (c) the elapsed (d) the height above the ground

OpenStudy (anonymous):

a) Here , at the top of the flight , the velocity is zero

OpenStudy (anonymous):

b) as the acceleration of the watermelon is same at all time in the air, acceleration = -g = -9.8 m.s^2

OpenStudy (anonymous):

c) time taken to reach top is t Using first equation of motion v = u + a*t 0 = -9.5 + 9.8 * t t = 0.97 s the time taken to reach the top is 0.97 s

OpenStudy (anonymous):

d) height above the ground = v^2/2g + Ho height above the ground = 9.5^2/*2 * 9.8 + 1.2 height above the ground = 5.81 m the height above the ground is 5.81 m

OpenStudy (ranya99):

Thanks a Lot